# SC-Diff: 3D Shape Completion with Latent Diffusion Models

### v1

@@ 3.2 Latent Diffusion Model for 3D Shapes
3.2.1 Forward Diffusion Process
The forward diffusion process is a fixed Markov Chain, adding Gaussian noise to a latent sample $\mathbf{z}$ over time steps $t\in{1,...,T}$, following a variance schedule $\beta_{1},\beta_{2},...\beta_{T}$ [15] (see Eq. 4, Fig. 2).
@@ 3.2.1 Forward Diffusion Process
$q(\mathbf{z}_{t}|\mathbf{z}_{t-1}):=\mathcal{N}(\mathbf{z}_{t};\sqrt{1-\beta_{t}}\mathbf{z}_{t-1},\beta_{t}\mathbf{I}).$ (4)
3.2.2 Reverse Diffusion Process
@@
$p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t}):=\mathcal{N}(\mathbf{z}_{t-1};\bm{\mu}_{\theta}(\mathbf{z}_{t},t,\mathbf{c},\mathbf{I}),\bm{\Sigma}_{\theta}(\mathbf{z}_{t},t,\mathbf{c},\mathbf{I}))$ (5)
Where $\mathbf{c}$ and $\mathbf{I}$ are the partial TSDF and RGB image respectively, and $\bm{\mu}_{\theta}$ and $\bm{\Sigma}_{\theta}$ are the mean and covariance predicted by our network in latent space. Learning the reverse diffusion process involves estimating these mean and covariance. However, it has been shown experimentally that setting $\bm{\Sigma}_{\theta}(\mathbf{z}_{t},t,\mathbf{c},\mathbf{I})=\sigma_{t}^{2}\mathbf{I}$ with $\sigma_{t}^{2}=\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ with $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$ provides good results, while removing the need to learn the covariance.
We use the reparametrization and objective proposed by  [15] (see Eq. 6) and focus on predicting the noise $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$ added during the forward process.
$L_{simple}(\theta):=\mathbb{E}_{t,\mathbf{z},\bm{\epsilon}}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c},\mathbf{I})\|_{2}^{2}\right].$ (6)
At training time, we take a complete TSDF voxel grid $\mathbf{X}$  and compute a latent code $\mathbf{z}_{0}$ using our VQ-VAE’s encoder $\mathcal{E}$, then we sample a noised latent code $\mathbf{z}_{t}$ at time $t$. The denoising diffusion model predicts the noise $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$ and is trained with the objective shown in Eq. 6.
In contrast, at inference time, we sample $\mathbf{\hat{z}}$ by gradually denoising a noise variable $\textbf{z}_{T}$ sampled from the standard normal distribution $\mathcal{N}(\mathbf{z}_{T};\mathbf{0},\mathbf{I})$ and then use our pretrained decoder $\mathcal{D}$ and codebook $\mathcal{Z}$ to map the denoised code $\mathbf{\hat{z}}$ back to a predicted complete TSDF voxel grid $\mathbf{\hat{X}}$.
Our denoising diffusion model $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$ is implemented as a time-conditional 3D U-Net  [15, 31].
@@ 3.3 Conditioning mechanisms
We propose the use of two conditioning mechanisms, as depicted in Fig. 2: cross-attention on CLIP features [27] from an object image and spatially consistent feature aggregation from a partial scan’s TSDF voxel grid.
3.3.1 Conditioning 3D shape generation on RGB Images
Diffusion models are capable of modeling conditional distributions of the form $p(\mathbf{z}|\mathbf{y})$, which allows us to implement a conditional denoising autoencoder $\epsilon_{\theta}(z_{t},t,y)$ where $y$ is a conditioning input.
@@ 3.3.1 Conditioning 3D shape generation on RGB Images
Inspired by latent diffusion models [31], we extend the U-Net backbone with a cross-attention mechanism. This allows learning the distribution of latent codes $p(\mathbf{z}|\mathcal{E}_{I}(\mathbf{I}))$ conditioned on pre-trained CLIP encoder features of an RGB image of the 3D object rendered from the pose where the partial scan was taken.
The CLIP encoder $\mathcal{E}_{I}$ projects the RGB Image $\mathbf{I}\in\mathbb{R}^{3\times 256\times 256}$ to an intermediate representation $\mathcal{E}_{I}(\mathbf{I})\in\mathbb{R}^{M\times d_{\mathcal{E}_{I}}}$ where $d_{\mathcal{E}_{I}}=768$ is the dimension of the CLIP features and $M=50$ corresponds to the flattened spatial dimension of CLIP’s encoding. This representation is then mapped to the flattened intermediate layers of the time-conditional UNet via cross-attention layers.
3.3.2 Conditioning 3D shape generation on Partial Scans
Simultaneously, to enhance control over the spatial composition and address limitations of image-only conditioning, we introduce a second mechanism inspired by ControlNet and DiffComplete [42, 8].
First, the corrupted ground truth latent code $\mathbf{z}_{t}$ and partial shape $\mathbf{c}$ are encoded into multi-resolution feature maps of the same size $\mathcal{E}_{\epsilon}(\mathbf{z}_{t})$ and $\mathcal{E}_{c}(\mathbf{c})$ respectively. While the diffusion model forwards $\mathcal{E}_{\epsilon}(\mathbf{z}_{t})$, the control branch propagates the fused features $\mathbf{f}=\mathcal{E}_{\epsilon}(\mathbf{z}_{t})+\mathcal{E}_{c}(\mathbf{c})$ into deeper layers. The control branch is connected to the diffusion model with projection 1x1 convolutional layers such that the feature maps have consistent sizes and can be aggregated. The features that enter the i-th decoder block of the diffusion model are denoted as:
$d^{(i)}=[\mathcal{D}_{z}^{(i-1)}(\mathbf{z}_{t}),F_{z}^{(i)}(\mathbf{z}_{t})+\phi^{(i)}(F_{c}^{(i)}(\mathbf{f}))]$ (7)
Where $\phi^{(i)}$ is the 1x1 convolutional projection layer, $F_{z}^{(i)}(\mathbf{z}_{t})$ are the encoder features of the diffusion model’s i-th encoder block that are passed through residual connections to the decoder block $\mathcal{D}_{z}^{(i-1)}$. $F_{c}^{(i)}(\mathbf{f})$ corresponds to the features of the control branch’s i-th encoder block, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the concatenation operation.
Combining the presented conditioning approaches allows the model to leverage global shape cues from images and detailed 3D information from partial scans, facilitating realistic and high-fidelity shape completion.
@@ 3.3.2 Conditioning 3D shape generation on Partial Scans
$l1$-err $\downarrow$ 3D-EPN AutoSDF PatchComplete DiffComplete Ours Ours Ours
@@ 4 Experiments
We evaluate our method in the task of shape completion on two large-scale shape completion benchmarks. The 3D-EPN [12] dataset contains 25,590 object instances of eight classes from ShapeNet [5]. For each instance, several partial scans of varying completeness are provided as $32^{3}$ TSDF grids, while complete shapes are represented as $32^{3}$ or $64^{3}$ truncated unsigned distance functions. The Patchcomplete [29] benchmark emphasizes completing objects from unseen categories. It contains 5000 partial/complete pairs from the synthetic Shapenet [5] and 11,000 partial/complete pairs taken from objects in the scenes of the real-world ScanNet dataset [9].
We first train a VQ-VAE for 150k steps. Next, we train our conditional shape completion network (diffusion model + control branch) using the same data split for other 200k steps, while the CLIP encoder remains frozen with pre-trained weights. In both cases we use Adam optimizer [43] with a learning rate of $1\times 10^{-4}$ and $2.5\times 10^{-5}$, respectively. We set our batch size to $8$ and trained both VQ-VAE and our shape completion network on a single Nvidia A40 GPU with 48GB of VRAM. While the VQ-VAE can be trained in 48 hours, training the diffusion model takes about 120 hours. On the other hand, at inference we use the formulation of DDIM [34], needing only 100 denoising steps which take about 15 seconds per sample.
@@ 4.1 Quantitative Evaluation
As commonly used in the field, we evaluate the $l_{1}$ error across all voxels for known object classes as well as the chamfer distance (CD), and the Intersection over Union (IoU) between predicted and ground truth shapes for unknown classes.
Figure 3: Shape completion results on the test set of the 3D-EPN [12] benchmark. DiffComplete [8] results are not shown due to the lack of open-source code and weights.
4.1.1 3D-EPN Benchmark - Known Object Classes:
Our method demonstrates competitive performance when directly compared to the state-of-the-art techniques by training class-specific models, outperforming them by a $\downarrow 12\%$ reduction in l1 (see Tab. 1). It is noteworthy that our class-agnostic model still has similar performance to the class-specific state-of-the-art model Diffcomplete [8], and outperforms the other compared class-agnostic models, which showcases its great generalization capabilities. Unlike the compared techniques, our method additionally uses the RGB image from the partial scans. Furthermore, it is important to highlight that the compared approaches are designed to operate on a $32^{3}$ resolution, whereas our model is capable of processing $64^{3}$ TSDF voxel grids. As shown qualitatively in Fig. 1, our diffusion-based approach excels in predicting fine-detailed shapes from the partial input scans.
@@
Bag 5.01 / 73.8 5.81 / 56.3 3.94 / 77.6 3.86 / 78.30 3.79 / 78.26
Lamp 8.07 / 47.2 6.57 / 39.1 4.68 / 56.4 4.80 / 57.90 4.74 / 59.95
Bathtub 4.21 / 57.9 5.17 / 41.0 3.78 / 66.30 3.52 / 68.90 3.67 / 65.88
Bed 5.84 / 58.4 6.01 / 44.6 4.49 / 66.8 4.16 / 67.10 4.40 / 67.05
Basket 7.90 / 54.0 6.70 / 39.8 5.15 / 61.0 4.94 / 65.50 4.89 / 68.45
Printer 5.15 /73.6 5.82 / 49.9 4.63 / 77.6 4.40 / 76.80 4.36 / 76.77
Laptop 3.90 / 62.0 4.81 / 51.1 3.77 / 63.8 3.52 / 67.40 3.41 / 68.35
Bench 4.54 / 48.3 4.31 / 39.5 3.70 / 53.9 3.56 / 58.20 3.39 / 61.13
Avg. 5.58 / 59.4 5.86 / 45.2 4.27 / 65.4 4.10 / 67.50 4.08 / 68.17
@@
Bag 8.83 / 53.7 9.30 / 48.7 8.23 / 58.3 7.05 / 48.5 7.41 / 50.02
Lamp 14.3 / 20.7 11.2 / 24.4 9.42 / 28.4 6.84 / 30.5 6.39 / 33.17
Bathtub 7.56 / 41.0 7.84 / 36.6 6.77 / 48.0 8.22 / 48.50 8.09 / 48.44
Bed 7.76 / 47.8 7.91 / 38.0 7.24 / 48.4 7.20 / 46.6 6.91 / 48.63
Basket 7.74 / 36.5 7.54 / 36.1 6.60 / 45.5 7.42 / 59.2 6.38 / 62.15
Printer 8.36 / 63.0 9.66 / 49.9 6.84 / 70.5 6.36 / 74.5 7.10 / 69.06
Avg. 9.09 / 44.0 8.90 / 38.9 7.52 / 49.5 7.18 / 51.3 7.04 / 51.91
Figure 4: Qualitative Results on the Patchcomplete [29] benchmark for unseen categories.
4.1.2 Patchcomplete Benchmark - Unseen Object Classes:
For object categories not observed during training our method outperforms competitors in several specific categories such as bag, basket, lamp, and bench, and it also shows a marginal overall improvement for both synthetic and real-world data (see Tab. 2). Such outcomes can be attributed to the efficacy of our dual conditioning mechanisms, which effectively guide the latent diffusion model’s generative process, thereby ensuring the completion of shapes with a high grade of fidelity. Fig. 4 shows some of the predictions performed by our method on this benchmark, showcasing its superior ability to complete shapes from unseen categories with inputs of various levels of completeness.
@@ 4.1.2 Patchcomplete Benchmark - Unseen Object Classes:
Figure 5: Our approach produces diverse but realistic and detailed multimodal shape completions for the same input partial scans.
4.1.3 Comparison to SDFusion [6]:
Metric SDFusion Ours
l1-error $\downarrow$ 0.086 0.068
CD $\downarrow$ 0.058 0.025
IoU $\uparrow$ 0.38 0.55
Table 3: Comparison between SDFusion[6] and our method.
Even though our purpose was to perform shape completion from partial scans, we also compared the performance of our model with a state of the art method for 3D Latent Diffusion: SDFusion [6]. For this, we adopted their formulation of shape completion where the partial TSDFs are obtained by masking out big regions of the ground truth shapes, and finetuned our model for $50k$ further steps.
As shown in Tab. 3. Our method significantly outperforms SDFusion [6] in l1-error, chamfer distance, and IoU. Similarly, Fig. 6 shows how our method produces both more plausible and more diverse shapes.
Figure 6: Shape completion comparison with SDFusion[6]. Our method produces more plausible and diverse results.
4.2 Ablation Studies
4.2.1 Influence of 2D Loss on VQ-VAE Performance:
Table 4: Effect of 2D losses on VQ-VAE’s reconstructions.
2D Loss weight $(\gamma)$ $l1$-err $\downarrow$ IoU $\uparrow$
@@ 4.2.1 Influence of 2D Loss on VQ-VAE Performance:
We extend the original VQ-VAE [36] by incorporating supplementary 2D losses that operate on 2D renderings of the reconstructed TSDFs. Tab. 4 demonstrates the beneficial impact of these added 2D objectives on the autoencoder’s shape reconstructions, evidencing a substantial enhancement in reconstruction quality, with a $\downarrow 54\%$ reduction in $l1$ error and a $\uparrow 6\%$ increase in IoU.
4.2.2 Effects of different conditioning Types:
@@
Image-based Partial TSDF Avg. $l1$-err $\downarrow$
@@
Tab. 5 shows the impact of implementing our conditioning mechanisms by evaluating performance on the 3DEPN benchmark under three distinct scenarios: image-based conditioning alone, partial shape-based conditioning alone, and the concurrent application of both conditioning strategies. While relying solely on image-based conditioning yields suboptimal performance attributable to the inherent uncertainties in the image-to-shape reconstruction task, the synergy of image-based and partial shape conditionings results in a notable $15\%$ reduction in $l1$-error respect to partial shape conditioning only. This demonstrates that while each conditioning mechanism contributes uniquely, their combination harnesses their respective strengths to significantly enhance the shape completion process, allowing for a competitive performance even when we train a single model to predict complete shapes for multiple object categories.

### v2

@@ 3.2 Latent Diffusion Model for 3D Shapes
The forward diffusion process is a fixed Markov chain, adding Gaussian noise to a latent sample $\mathbf{z}$ over time steps $t\in{1,...,T}$, following a variance schedule $\beta_{1},\beta_{2},...,\beta_{T}$ [18].
@@
$q(\mathbf{z}_{t}|\mathbf{z}_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}\mathbf{z}_{t-1},\beta_{t}\mathbf{I}).$ (4)
Importantly, the forward process allows to sample $\mathbf{z}_{t}$ at an arbitrary timestep $t$ in closed form:
$q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0},(1-\bar{\alpha}_{t})\mathbf{I})\text{,}$ (5)
where $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$. This formulation makes it possible to sample arbitrary $\mathbf{z}_{t}$ without having to process $\mathbf{z}_{0}...,\mathbf{z}_{t-1}$.
@@
$p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t}):=\mathcal{N}(\bm{\mu}_{\theta}(\mathbf{z}_{t},t,\mathbf{c},\mathbf{I}),\bm{\Sigma}_{\theta}(\mathbf{z}_{t},t,\mathbf{c},\mathbf{I}))\text{,}$ (6)
where $\mathbf{c}$ and $\mathbf{I}$ denote the partial TSDF and RGB image, respectively, and $\bm{\mu}_{\theta}$ and $\bm{\Sigma}_{\theta}$ denote the mean and covariance predicted by our network in latent space. Learning the reverse diffusion process involves estimating $\bm{\mu}_{\theta}$ and $\bm{\Sigma}_{\theta}$. However, it has been shown experimentally that setting $\bm{\Sigma}_{\theta}(\mathbf{z}_{t},t,\mathbf{c},\mathbf{I})=\sigma_{t}^{2}\mathbf{I}$ with $\sigma_{t}^{2}=\tilde{\beta}_{t}=(1-\bar{\alpha}_{t-1})/({1-\bar{\alpha}_{t}})\beta_{t}$, $\alpha_{t}:=1-\beta_{t}$, and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$ provides good results, while removing the need to learn the covariance.
We use the reparametrization and objective $\mathcal{L}_{DDPM}$ proposed by [18] (see Eq. 7) and focus on predicting the noise $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$ added during the forward process.
$\mathcal{L}_{DDPM}:=\mathbb{E}_{t,\mathbf{z},\bm{\epsilon}}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c},\mathbf{I})\|_{2}^{2}\right].$ (7)
At training time, we take a complete TSDF voxel grid $\mathbf{X}$  and compute a latent code $\mathbf{z}_{0}$ using our VQ-VAE’s encoder $\mathcal{E}$. Then, we sample a noised latent code $\mathbf{z}_{t}$ at time $t$. The denoising diffusion model predicts the noise $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$ and is trained with the objective shown in Eq. 7. In contrast, at inference time, we sample $\mathbf{\hat{z}}$ by gradually denoising a noise variable $\mathbf{z}_{T}$ sampled from the standard normal distribution $\mathcal{N}(\mathbf{z}_{T};\mathbf{0},\mathbf{I})$. We then use our pretrained decoder $\mathcal{D}$ and codebook $\mathcal{Z}$ to map the denoised code $\mathbf{\hat{z}}$ back to a predicted complete TSDF voxel grid $\mathbf{\hat{X}}$. Our denoising diffusion model $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$ is implemented as a time-conditional 3D U-Net  [18, 35].
@@ 3.3 Conditioning Mechanisms
We propose the use of two complementary conditioning mechanisms, as illustrated in Fig. 2: cross-attention on CLIP features [30] extracted from an object image, and spatially consistent feature aggregation from the TSDF voxel grid of a partial scan.
@@
Conditioning on RGB Images Diffusion models can model conditional distributions of the form $p(\mathbf{z}|\mathbf{y})$, which allows us to implement a conditional denoising autoencoder $\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$, where $\mathbf{y}$ denotes a conditioning input. Following the trend of latent diffusion models [35], we extend the U-Net backbone with a cross-attention mechanism. This enables learning the distribution of latent codes $p(\mathbf{z}|\mathcal{E}_{I}(\mathbf{I}))$ conditioned on pre-trained CLIP encoder features of an RGB image $\mathbf{I}$ of the 3D object, rendered from the viewpoint corresponding to the partial scan. The CLIP encoder $\mathcal{E}_{I}$ projects the RGB image $\mathbf{I}\in\mathbb{R}^{3\times 256\times 256}$ to an intermediate representation $\mathcal{E}_{I}(\mathbf{I})\in\mathbb{R}^{M\times d_{\mathcal{E}_{I}}}$, where $d_{\mathcal{E}_{I}}=768$ is the feature dimension and $M=50$ corresponds to the flattened spatial dimension of the CLIP encoding. This representation is then injected into the intermediate layers of the time-conditional U-Net via cross-attention layers.
Conditioning on Partial Scans To improve control over spatial structure and overcome the limitations of image-only conditioning, we introduce a second mechanism inspired by ControlNet and DiffComplete [46, 9]. Instead of artificially corrupting the ground-truth TSDF map, we employ a mapping pipeline to obtain realistic partial reconstructions from a single viewpoint [27, 10]. Details of the partial TSDF preprocessing are provided in the supplementary. The corrupted ground-truth latent code $\mathbf{z}_{t}$ and the partial shape $\mathbf{c}$ are encoded into multi-resolution feature maps of the same size, $\mathcal{E}_{\epsilon}(\mathbf{z}_{t})$ and $\mathcal{E}_{c}(\mathbf{c})$, respectively. During the forward pass of the diffusion model, $\mathcal{E}_{\epsilon}(\mathbf{z}_{t})$ is propagated through the U-Net, while the control branch fuses the features as $\mathbf{f}=\mathcal{E}_{\epsilon}(\mathbf{z}_{t})+\mathcal{E}_{c}(\mathbf{c})$ and forwards them into deeper layers. The control branch connects to the diffusion model via $1\times 1$ convolutional layers to ensure consistent feature map dimensions for aggregation. The features entering the $i$-th decoder block of the diffusion model are expressed as:
$d^{(i)}=[\mathcal{D}_{z}^{(i-1)}(\mathbf{z}_{t}),F_{z}^{(i)}(\mathbf{z}_{t})+\phi^{(i)}(F_{c}^{(i)}(\mathbf{f}))]\text{,}$ (8)
where $\phi^{(i)}$ is the $1\times 1$ convolutional projection layer, $F_{z}^{(i)}(\mathbf{z}_{t})$ are the encoder features of the diffusion model’s $i$-th encoder block passed via residual connections to the decoder block $\mathcal{D}_{z}^{(i-1)}$, and $F_{c}^{(i)}(\mathbf{f})$ are the features from the control branch’s $i$-th encoder block. Here, $\cdot$$\cdot$] denotes concatenation. This design ensures spatially consistent integration of partial scan features into the reverse diffusion process. Combining both conditioning approaches allows the model to leverage global shape cues from images while incorporating detailed 3D information from partial scans, resulting in realistic and high-fidelity shape completion.
@@
$l_{1}$ error $\downarrow$ 3D-EPN AutoSDF PatchComplete DiffComplete Ours Ours Ours
@@ 4 Experiments
We evaluate our method in the task of shape completion on two large-scale shape completion benchmarks. The 3D-EPN [13] dataset contains 25,590 object instances of eight classes from ShapeNet [6]. For each instance, several partial scans of varying completeness are provided as $32^{3}$ TSDF grids, while complete shapes are represented as $32^{3}$ or $64^{3}$ truncated unsigned distance functions. The PatchComplete [32] benchmark emphasizes completing objects from unseen categories. It contains 5,000 partial/complete pairs from the synthetic ShapeNet [6] and 11,000 partial/complete pairs from objects in the scenes of the real-world ScanNet dataset [12].
We first train a VQ-VAE for 150,000 steps. Next, we train our conditional shape completion network (diffusion model and control branch) using the same data split for other 200,000 steps, while the CLIP encoder remains frozen with pre-trained weights. In both cases we use the Adam optimizer [47] with a learning rate of $1\times 10^{-4}$ and $2.5\times 10^{-5}$, respectively. We set the batch size to $8$ and trained both the VQ-VAE and the shape completion network on a single NVIDIA A40 GPU with 48 GB of VRAM. While the VQ-VAE can be trained in 48 hours, training the diffusion model takes about 120 hours. On the other hand. At inference, we use the DDIM formulation [38], requiring only 100 denoising steps, which take approximately 15 seconds per sample.
@@ 4.1 Quantitative Evaluation
As commonly used in the field, we evaluate the $l_{1}$ error across all voxels for known object classes as well as the chamfer distance (CD), and the intersection over union (IoU) between predicted and ground-truth shapes for unknown classes.
[…]
Figure 3: Qualitative results on the 3D-EPN [13] test set. DiffComplete [9] results are excluded due to unavailable open-source weights. Leveraging improved model scalability, our method reconstructs finer geometric details while maintaining faithful to the input shapes.
3D-EPN Benchmark – Known Object Classes Our method achieves superior performance compared to state-of-the-art approaches. On the standard $32^{3}$ resolution benchmark, our class-agnostic model performs on par with the class-specific DiffComplete [9], while widely surpassing all other class-agnostic baselines by $47\%$. This demonstrates strong generalization across categories without relying on class-specific supervision. Moreover, when trained in a class-specific setting, our model outperforms DiffComplete by a $12\%$ reduction in $l_{1}$ error while being $30\%$ more memory-efficient, highlighting the effectiveness of our latent formulation and multimodal conditioning. Importantly, unlike prior methods limited to $32^{3}$ due to little or no compression, our approach scales to $64^{3}$ TSDF voxel grids while remaining efficient, enabling higher-fidelity completions with finer geometric details. As shown in Tab. 5, we further analyze the impact of different conditioning mechanisms. Notably, even when conditioned only on the partial TSDF input, our class-agnostic model still outperforms the state-of-the-art by a $19\%$ reduction in $l_{1}$ error, underlining the strength of our latent diffusion framework. Incorporating RGB guidance further enhances reconstruction quality, as illustrated by the qualitative results in Fig. 1, where our model excels at recovering realistic and detailed shapes from incomplete observations.
@@
Bag 5.01 / 73.8 5.81 / 56.3 3.94 / 77.6 3.86 / 78.3 3.79 / 78.3
Lamp 8.07 / 47.2 6.57 / 39.1 4.68 / 56.4 4.80 / 57.9 4.74 / 60.0
Bathtub 4.21 / 57.9 5.17 / 41.0 3.78 / 66.3 3.52 / 68.9 3.67 / 65.9
Bed 5.84 / 58.4 6.01 / 44.6 4.49 / 66.8 4.16 / 67.1 4.40 / 67.1
Basket 7.90 / 54.0 6.70 / 39.8 5.15 / 61.0 4.94 / 65.5 4.89 / 68.5
Printer 5.15 / 73.6 5.82 / 49.9 4.63 / 77.6 4.40 / 76.8 4.36 / 76.8
Laptop 3.90 / 62.0 4.81 / 51.1 3.77 / 63.8 3.52 / 67.4 3.41 / 68.4
Bench 4.54 / 48.3 4.31 / 39.5 3.70 / 53.9 3.56 / 58.2 3.39 / 61.1
Avg. 5.58 / 59.4 5.86 / 45.2 4.27 / 65.4 4.10 / 67.5 4.08 / 68.2
@@
Bag 8.83 / 53.7 9.30 / 48.7 8.23 / 58.3 7.05 / 48.5 7.41 / 50.0
Lamp 14.3 / 20.7 11.2 / 24.4 9.42 / 28.4 6.84 / 30.5 6.39 / 33.2
Bathtub 7.56 / 41.0 7.84 / 36.6 6.77 / 48.0 8.22 / 48.5 8.09 / 48.5
Bed 7.76 / 47.8 7.91 / 38.0 7.24 / 48.4 7.20 / 46.6 6.91 / 48.6
Basket 7.74 / 36.5 7.54 / 36.1 6.60 / 45.5 7.42 / 59.2 6.38 / 62.2
Printer 8.36 / 63.0 9.66 / 49.9 6.84 / 70.5 6.36 / 74.5 7.10 / 69.1
Avg. 9.09 / 44.0 8.90 / 38.9 7.52 / 49.5 7.18 / 51.3 7.04 / 51.9
[…]
Figure 4: Qualitative results on the PatchComplete [32] benchmark for unseen categories. Note the generally smoother surfaces and sharper edges produced by our model.
PatchComplete Benchmark – Unseen Object Classes For object categories not observed during training, our method outperforms competitors in almost all specific categories, demonstrating an improvement for both synthetic and real-world data (see Tab. 2). Fig. 4 shows some of the predictions performed by our method on this benchmark, showcasing its superior ability to complete shapes from unseen categories with inputs of various levels of completeness.
@@
[…]
Figure 5: Our approach produces diverse but realistic and detailed multifaceted shape completions for the same input partial scans, for both real-world (top) and synthetic (bottom) examples.
Metric SDFusion [7] ShapeFormer [42] Ours
$l_{1}$ error $\downarrow$ 0.086 0.097 0.068
CD $\downarrow$ 0.058 0.073 0.025
IoU $\uparrow$ 0.38 0.31 0.55
Table 3: Quantitative comparison between SDFusion[7], ShapeFormer [42], and our method on the ShapeNet dataset [6].
Comparison to Non-TSDF-Based Approaches While our method and other previously discussed approaches operate on TSDF grids, the methods compared in Tab. 3 use different 3D representations: SDFusion [7] relies on signed distance functions (SDFs), and ShapeFormer [42] operates on point clouds. For a fair comparison with SDFusion [7], we followed their shape completion strategy by generating partial inputs through masking significant portions of the ground-truth shapes and finetuned our model for an additional 50,000 steps. The results in Tab. 3 show that our TSDF-based approach substantially outperforms both SDFusion [7] and ShapeFormer [42] across all metrics. Qualitative examples in Fig. 5 further illustrate that our method produces more plausible shapes with higher structural fidelity and greater shape diversity. Additional qualitative comparisons can be found in the supplementary.
Table 4: Effect of 2D losses on the VQ-VAE reconstructions.
2D loss weight $(\gamma_{A})$ $l_{1}$ error $\downarrow$ IoU $\uparrow$
@@
Image-based Partial TSDF Avg. $l_{1}$ error $\downarrow$
@@
4.2 Ablation Studies
Effect of Increasing Resolutions Fig. 6 illustrates shape completions at different output resolutions. While our low-resolution model achieves state-of-the-art performance, the reduced grid size inevitably oversmoothes complex geometries, such as shelves and chairs. In contrast, our higher-resolution outputs at $64^{3}$ preserve significantly finer structural details.
[…]
Figure 6: Completions for various resolutions on ShapeNet [6].
Influence of 2D Loss on VQ-VAE Performance We extend the original VQ-VAE [40] by incorporating supplementary 2D losses that operate on 2D renderings of the reconstructed TSDFs. Tab. 4 demonstrates the beneficial impact of these added 2D objectives on the autoencoder’s shape reconstructions, showing a substantial enhancement in reconstruction quality, with a $54\%$ reduction in $l_{1}$ error and a $6\%$ increase in IoU. Qualitative illustrations of the efficacy of integrating a 2D objective to create high-fidelity reconstructions are shown in the appendix. As shown, our training strategy results in reconstructions that more accurately reflect the original input shapes. Such high-fidelity reconstruction is essential, given that the efficient encoding of the TSDF space into a compact latent representation is critical for the success of our conditional shape completion diffusion model.
Effects of Different Conditioning Types Tab. 5 evaluates the contribution of our conditioning mechanisms on the 3D-EPN benchmark across three scenarios: image-based conditioning alone, partial shape-based conditioning alone, and their combination. While each conditioning mechanism provides unique benefits, their concurrent use leads to a further $15\%$ reduction in $l_{1}$ error compared to using partial shape conditioning alone. This highlights that the combination effectively leverages complementary information and enables a single model to perform competitive shape completion across multiple object categories.